OpenStudy (kramse18):

What are the first 3 non-zero terms in the power series of arcsin(2x)?

OpenStudy (bobo-i-bo):

In general, do you know how to find the terms of a power series?

OpenStudy (kramse18):

sort of

OpenStudy (bobo-i-bo):

Go on :)

OpenStudy (kramse18):

I was getting 1+(1/2)arcsin(4)+(1/2)arcsin(192) but you can't take the arcsin of 4

OpenStudy (bobo-i-bo):

That's because you're applying the formula wrong... the power series is a function of x. Yet your sum is not a function of x.

OpenStudy (kramse18):

So then what do I do?

OpenStudy (bobo-i-bo):

Given f(x), the power series of f(x) is (assuming we are doing the power series around 0): \[f(x)=\sum_{n=0}^{\infty}\frac{ 1 }{ n! } f^{(n)}(0)x^n\]

OpenStudy (bobo-i-bo):

In this case f(x)=arcsin(2x). Then you apply the power series formula above. Understanding me? :P

OpenStudy (kramse18):

so arcsin(2x) would go where the f^n is?

OpenStudy (bobo-i-bo):

Oh sorry, the notation \[f^{(n)}(x)\] is another notation for: \[\frac{ d^ny }{ dx^n }(x)\]

OpenStudy (bobo-i-bo):

Let me give you an example.

OpenStudy (bobo-i-bo):

If f(x)=sin(x). Then the power series of f(x) i.e. the power series of sin(x) is: \[f(x)=\sum_{n=0}^{\infty}\frac{1}{n!}\frac{d^ny}{dx^n}(0)x^n=\frac{1}{0!}y(0)x^0+\frac{1}{1!}\frac{dy}{dx}(0)x^1+\frac{1}{2!}\frac{d^2y}{dx^2}(0)x^2+\frac{1}{3!}\frac{d^3y}{dx^3}(0)x^3...\] =\[\sin(0)x^0+(\frac{d}{dx}(\sin(0)))x^1+\frac{1}{2}(\frac{d^2}{dx^2}(\sin(0)))x^2+\frac{1}{6}(\frac{d^3}{dx^3}(\sin(0)))x^3+...\]

OpenStudy (bobo-i-bo):

\[=\sin(0)x^0+\cos(0)x^1-\sin(0)x^2-\frac{1}{6}\cos(0)x^3+...\]

OpenStudy (bobo-i-bo):

Lol, it looks really messy... but do you get what to do now?

OpenStudy (bobo-i-bo):

Do you follow what i've done?

OpenStudy (kramse18):

I think I do. Thanks

OpenStudy (bobo-i-bo):

np :)